OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

wher em = slope and \[x_1,y_1\] is a point on that line

OpenStudy (anonymous):

so u can count c and d out

OpenStudy (anonymous):

correct just substitute th givens into my equation

OpenStudy (anonymous):

b

OpenStudy (anonymous):

\[y-6=-3(x-(-1))\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):